x/(x^2-9)+4/(x+3)=3/(x^2-9)
Answer by stanbon(75887)

x/(x^2-9)+4/(x+3)=3/(x^2-9)
----
Note: x^2-9 = (x-3)(x+3)
----------------
Multiply thru by x^2-9 to get:
x + 4(x-3) = 3
---
5x-12 = 3
5x = 15
x = 3
============
Cheers,
Stan H.
==========